On January 1, 2021, the Brunswick Hat Company adopted the dollar-value LIFO retail method. The following data are available for 2021:
 


CostRetail
Beginning inventory$72,900$135,000
Net purchases
102,150
232,000
Net markups


5,000
Net markdowns


10,000
Net sales


200,000
Retail price index, 12/31/2021


1.08


Required:
Calculate the estimated ending inventory and cost of goods sold for 2021 using the information provided. (Do not round intermediate calculations.)
Answer #1

step 1)


costretail
begin inventory72,900135,000
net purchase102,150232,000
net markup
5,000
net markdown                                   
(10,000)
purchases175,050362,000
net sales
(200,000)
end inventory at retail
162,000
end inventory at cost
(80,190)
Cost of goods sold94,860

72,900/135,000= 0.54 or 54%

step 1) 232,000-10,000+5,000=227,000

102,150/227,000= 0.45 or 45%


Step 2) 135,000*54%*1.00= 72,900


step 3) 162,000/1.08= 150,000-135,000= 15,000

15,000*45%*1.08= 7,290

72,900+7,290= 80,190


Answer:

ending inventory at retail= 162,000

ending inventory at cost= 80,190

cost of goods sold= 94,860

answered by: anonymous